OpenStudy (anonymous):

f(x)=2x^2-x ;g(x)=x^(-1/2) f(g(x))

OpenStudy (anonymous):

f(g(x))=f(x^(1/2))=2( x^(1/2) )^2 - x^(1/2)= 2x -x^(1/2)

OpenStudy (anonymous):

sorry forgot to add the neg. (^-1/2)

OpenStudy (mathstudent55):